News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 25 January 2002 10.00am EST

During the last week the observing schedule was interrupted when the spacecraft transitioned to bright star hold following a long maneuver at 10.53pm EST on Jan 23. All actions were nominal and analysis of the real-time and dump data determined that the transition was due to a maneuver error of 150 arcsec exceeding the aspect camera star search box size of 120 arcsec. The error resulted from the combination of the specific eigenaxis alignment for the long maneuver and a slightly inaccurate on-board gyro scale-factor and alignment matrix. An updated gyro scale-factor and alignment matrix is currently under test and will be uplinked following approval. In the meantime, loads will be screened for similar maneuver cases.

A nominal recovery from bright star hold was completed at 2.40pm EST on Jan 24 and loads were rebuilt and resumed by 4pm in time for an observation of V471 Tau coordinated with HST. Observations of Cas A and IC342 were impacted and will be rescheduled in a future load.

Of note last week was the release of the Chandra cycle 4 NRA by NASA HQ. The proposal deadline is March 15, 2002.

The observing schedule for next week is shown below.

------------------------------------------
Q1458+718               ACIS-S       Jan 27 
CYG X-1                 ACIS-S/HETG  Jan 28 
Vela PWN                ACIS-S
Radiation Belts
SDSSP J130608+035626(3) ACIS-S       Jan 29 
NGC 3245                ACIS-S
NGC 4438                ACIS-S
CYG X-1                 ACIS-S
HS0248+3402             ACIS-S
0229+131                ACIS-S       Jan 30 
PSS 0311+0407           ACIS-S
G166.0+4.2              ACIS-I
HS0810+2554             ACIS-S
IRAS 05189-2524         ACIS-S
VLA J0636+1838          ACIS-I
Radiation Belts
Beta Cru                ACIS-S/HETG  Jan 31 
YY MEN, HD 32918        ACIS-S/HETG  Feb 1  
Radiation Belts                      Feb 2  
GX 301-2                ACIS-S/HETG  Feb 3

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports